OpenStudy (anonymous):

after driving 120km,4/7 of a trip is completed. How long is the total trip? How many more km must be driven to complete the tip?

OpenStudy (anonymous):

X is the length of the trip. X(4/7) = 120km is what you know. To find out the total trip, multiply 120km by the inverse of 4/7, which is 7/4, to get x by itself, leaving you with the length of the trip. Simply take this number and subtract it by the length traveled (120km) and you will have your answer for km remaining.

OpenStudy (anonymous):

To simplify = X(4/7) = 120km X (4/7) (7/4) = 120km (7/4) X = 210km (Length of the trip) 210km - 120km = 90km (length remaining)

OpenStudy (anonymous):

the entire trip is?

OpenStudy (anonymous):

Yes, refer to the work above. X = 210km, which is the entire length.